OpenStudy (anonymous):

Graph the linear equation. Find three points that solve the equation, then plot on the graph -2x-3y=-7

OpenStudy (anonymous):

3y=7-2x \[y=\frac{ 7-2x }{ 3 }\] put x=-1, \[y=\frac{ 7-2(-1) }{ 3 }=\frac{ 9 }{ 3 }=3\] the idea is to find the value of x and y integers. add 3 successively to find the value of y x=-1+3=2 x=2+3=5 find the corresponding values of y then make a graph.

OpenStudy (anonymous):

i have the graphs find the coordinates is toughest part for me

OpenStudy (anonymous):

@kool_22

OpenStudy (anonymous):

Deez nuts...

OpenStudy (loser66):

want to play with it?? I am in your side. Play??

OpenStudy (anonymous):

one point i have found namely (-1,3) find two other points similarly (2,_),(5,_) etc

OpenStudy (anonymous):

how would i find other two points

OpenStudy (anonymous):

@MrNood

OpenStudy (anonymous):

put x=2 in the above relation and find y similarly put x=5 and find y